### v1

# Rethinking Pan-sharpening: Principled Design, Unified Training, and a Universal Loss Surpass Brute-Force Scaling

Abstract
The field of pan-sharpening has recently seen a trend towards increasingly large and complex models, often trained on single, specific satellite datasets. This approach, however, leads to high computational overhead and poor generalization on full resolution data, a paradigm we challenge in this paper. In response to this issue, we propose PanTiny, a lightweight, single-step pan-sharpening framework designed for both efficiency and robust performance. More critically, we introduce multiple-in-one training paradigm, where a single, compact model is trained simultaneously on three distinct satellite datasets (WV2, WV3, and GF2) with different resolution and spectral information. Our experiments show that this unified training strategy not only simplifies deployment but also significantly boosts generalization on full-resolution data. Further, we introduce a universally powerful composite loss function that elevates the performance of almost all of models for pan-sharpening, pushing state-of-the-art metrics into a new era. Our PanTiny model, benefiting from these innovations, achieves a superior performance-to-efficiency balance, outperforming most larger, specialized models. Through extensive ablation studies, we validate that principled engineering in model design, training paradigms, and loss functions can surpass brute-force scaling. Our work advocates for a community-wide shift towards creating efficient, generalizable, and data-conscious models for pan-sharpening. The code is available at https://github.com/Zirconium233/PanTiny.
Figure 1: Our proposed PanTiny framework enables training a single, unified model on multiple datasets (WV2, WV3, GF2) simultaneously. This all-in-one approach achieves SOTA performance while maintaining a significantly smaller model size and lower computational cost compared to methods that require separate, specialized models for each dataset.
1 Introduction
Pan-sharpening, a fundamental image fusion task in remote sensing, aims to merge a high-resolution panchromatic (PAN) image with a lower-resolution multispectral (LRMS) image to generate a single high-resolution multispectral (HRMS) image. This fused image is crucial for numerous downstream applications, including land-cover classification, environmental monitoring, and urban planning (Masi et al. 2016; Yang et al. 2017). Early approaches were dominated by traditional methods such as Component Substitution (CS) (Carper et al. 1990; Chavez and Kwarteng 1989) and Multi-Resolution Analysis (MRA) (King and Wang 2001; Liu 2000), which, while efficient, often introduced spectral and spatial distortions due to their handcrafted prior. The advent of deep learning, particularly with models like PNN (Masi et al. 2016) and PanNet (Yang et al. 2017), revolutionized the field by learning complex mappings directly from data and significantly improving fusion quality. However, the recent pursuit of higher performance has led to a problematic trend. State-of-the-art (SOTA) methods increasingly rely on massive, complex architectures. For instance, CFDCNet (Li et al. 2025) achieves high metrics but at the cost of an astounding 55G FLOPs under 128x128 resolution images. Other models like Pan-Mamba (He et al. 2025) perform well but show limited generalization capabilities. Methods such as PanFlow (Yang et al. 2023) and diffusion-based models (Zhong et al. 2025) are often multi-step, complicating the inference process. Conversely, lightweight solutions like Pan-LUT (Cai et al. 2025), while fast, exhibit a noticeable performance gap. These specific shortcomings point to deeper, community-wide issues rooted in the pursuit of benchmark leadership over practical utility. A prevailing belief in "brute-force scaling" has led to models that are excessively large for the domain’s relatively small datasets (Deng et al. 2022). This issue is worsened by the community standard of training separate models for each satellite—a "one-dataset, one-model" philosophy that hinders both deployment efficiency and generalization. This reveals a critical issue: the majority of training for single-dataset models may contribute little to true, transferable generalization. This is evidenced by a finding in our appendix: a model fully converged on a source dataset shows nearly the same cross-domain performance as a simple baseline trained on that same source dataset for just one epoch. Secondly, the community standard of training separate models for each dataset (‘one-dataset, one-model’) hinders convenience and generalization. This is critical because true generalization is not merely about testing a single-dataset model on other datasets; as our appendix shows, such models are often just overfitting, with their cross-domain performance being matched by a model trained for just one epoch. In this work, we propose a comprehensive solution that challenges these norms. We first introduce PanTiny, a lightweight model that, AS illustrated in Figure 1 strikes a good balance between performance and efficiency, . Our extensive architectural ablations show that while scaling our model can further boost performance, it comes with diminishing returns, justifying our focus on efficiency. Second, we introduce a new “all-in-one” training paradigm in pan-sharpening domain, Through extensive experiments, we demonstrate that this approach not only simplifies the deployment pipeline but also significantly improves generalization on full-resolution data. We believe our findings support a shift towards unified models. This direction is inspired by the foundation model trend in other vision fields, suggesting a promising path for creating more generalizable pan-sharpening solutions. Finally, we design a powerful and universal composite loss function that significantly elevates the performance of all tested models, pushing the SOTA for metrics like GF2 PSNR into the 48-49 dB era. Our contributions are threefold:
  • We propose and validate a new “all-in-one” training paradigm in pan-sharpening domain, demonstrating its ability to improve generalization on full-resolution data across various models.
  • We propose PanTiny, a lightweight yet powerful model that strikes a good balance between performance and efficiency, whose design is methodically guided by key insights distilled from our extensive architectural ablations.
  • We introduce a universally effective composite loss function that consistently boosts the performance of diverse models, setting a new benchmark for the pan-sharpening field.
2 Related Work
2.1 Traditional Pan-sharpening
Traditional pan-sharpening methods are generally categorized into Component Substitution (CS), Multi-Resolution Analysis (MRA), and hybrid approaches. CS-based methods, such as Intensity-Hue-Saturation (IHS) (Carper et al. 1990) and Principal Component Analysis (PCA) (Chavez and Kwarteng 1989), project the MS image into a different space, replace one component with the PAN image, and then perform an inverse transform. These methods excel at enhancing spatial details but often suffer from significant spectral distortion. MRA-based methods, like those using wavelet transforms (King and Wang 2001) or smoothing filters (e.g., SFIM (Liu 2000)), decompose the images into different frequency bands and inject the high-frequency details from the PAN image into the MS image. MRA methods generally preserve spectral information better but can introduce spatial artifacts.
2.2 Deep Learning-based Pan-sharpening
The success of deep learning in computer vision spurred its application in pan-sharpening. PNN (Masi et al. 2016) was a pioneering work that used a simple three-layer CNN to learn the mapping from up-sampled MS and PAN images to the high-resolution MS output. PanNet (Yang et al. 2017) improved upon this by working in the high-frequency domain and introducing a spectral loss to better preserve color information. Subsequent works explored more complex CNN architectures, such as MSDCNN (Yuan et al. 2018), which used multi-scale features to improve fusion quality. These methods consistently outperformed traditional techniques, setting a new standard for the field.
2.3 Recent Advances and SOTA Models
The current landscape of pan-sharpening is dominated by advanced deep learning architectures. Inspired by successes in other vision tasks, researchers have incorporated Transformers (Zhou, Liu, and Wang 2022), State-Space Models (SSMs) like Mamba (He et al. 2025), and flow-based models (Yang et al. 2023). For instance, Pan-Mamba (He et al. 2025) leverages the efficiency of SSMs to achieve impressive results. The very recent CFDCNet (Li et al. 2025) has pushed performance metrics to new heights, but at the cost of an enormous computational load (55G FLOPs). Other approaches, such as PSCINN (Wang et al. 2024), utilize invertible neural networks to model the fusion process. While powerful, these models often come with a substantial increase in parameters and complexity. Furthermore, a common thread among these SOTA methods is their training protocol: they are almost exclusively trained and tested on a single dataset. Some works explore generalization by training on one dataset and testing on others (Chen et al. 2022), but the performance drop is often significant. The concept of an “all-in-one” model, trained jointly on multiple datasets for a single task, remains largely unexplored in pan-sharpening, representing a key opportunity that our work addresses.
3 Methodology
Our proposed method, PanTiny, is built on the principles of efficiency, simplicity, and empirical validation. We deliberately avoid overly complex operators and instead focus on a clean, effective architecture where each component’s inclusion is justified by extensive experiments. The overall architecture, shown in Figure 2, features a single-encoder design, a Transformer-based body for feature processing, and a simple convolutional refinement head.
Figure 2: The overall architecture of our proposed PanTiny framework. It consists of a single lightweight convolutional encoder for the MS input, a simple yet effective fusion module to integrate PAN information, a body of standard Transformer blocks for deep feature interaction, and a final convolutional layer for refinement.
3.1 Overall Architecture
Single Encoder
Unlike many methods that use separate encoders for PAN and MS inputs, we adopt an efficient single-encoder architecture. The upsampled MS image is first passed through a lightweight convolutional block to extract initial features. The PAN image is then integrated directly in the feature space via our fusion module. This design is highly parameter-efficient and, as our experiments show, forms the basis of a powerful and generalizable model.
Feature Fusion and Processing
Our investigation into fusion mechanisms revealed a surprising insight: in the multi-dataset training context, simplicity triumphs over complexity. We found that complex fusion strategies like cross-attention or the multi-layer “deepfusion” block from (He et al. 2025) actually degraded performance compared to a simple baseline. We attribute this to overfitting. Complex fusion modules tend to memorize dataset-specific artifacts. This "specialized knowledge" fails to generalize when the model is required to perform across multiple datasets, whereas a simpler module is forced to learn more robust, common features. The core of our network is a series of standard Transformer blocks, which effectively model long-range dependencies and perform deep feature interaction. After the Transformer body, we use a simple fusion block composed of two consecutive 3x3 convolutional layers (‘Enhanced Conv’). This design choice, validated in Table 7, proved to be the most effective and robust across all datasets.
Refinement Module
For the final reconstruction, we employ a single convolutional layer to map the fused features back to the desired high-resolution MS image. Our experiments (Table 8) confirmed that more elaborate refinement modules, such as those incorporating residual blocks or attention, offered no significant benefit and unnecessarily increased model size.
3.2 Transformer Block
While the overall structure is inspired by the original Transformer (Vaswani et al. 2017), our implementation uses a Pre-LayerNorm (Pre-LN) configuration for improved training stability. For an input feature map $X_{l-1}$, the output $X_{l}$ of a single Transformer block is computed as:
$X^{\prime}_{l}$ $=\text{CA}(\text{LN}(X_{l-1}))+X_{l-1}$ (1)
$X_{l}$ $=\text{GDFN}(\text{LN}(X^{\prime}_{l}))+X^{\prime}_{l}$ (2)
where LN denotes Layer Normalization, CA is our Channel Attention module, and GDFN is a Gated-DConv Feed-Forward Network.
Channel Attention (CA).
The CA module captures global context by performing self-attention across channel dimensions. Given an input $X\in\mathbb{R}^{B\times C\times H\times W}$, we first generate the query ($Q$), key ($K$), and value ($V$) projections via depth-wise convolutions. The attention map is then computed as:
$\text{Attention}(Q,K,V)=\text{Softmax}((Q_{n}K_{n}^{T})\cdot\tau)V_{n}$ (3)
where $Q_{n}$ and $K_{n}$ are L2-normalized query and key tensors, and $\tau$ is a learnable temperature parameter that scales the attention map. This design avoids the standard scaling by feature dimension, instead allowing the network to learn the optimal attention scaling.
Gated-DConv Feed-Forward Network (GDFN).
To enhance feature representation efficiently, we employ a gated feed-forward network. An input tensor is first projected to a higher-dimensional space and then split into two parallel paths, $X_{1}$ and $X_{2}$. The output is computed as:
$\text{GDFN}(X)=\text{Conv}_{\text{out}}(\text{GELU}(\text{DWConv}(X_{1}))\odot\text{DWConv}(X_{2}))$ (4)
where $\odot$ denotes element-wise multiplication. This gating mechanism allows for more dynamic and expressive feature transformations compared to a standard FFN.
Table 1: Main quantitative comparison. All models are trained simultaneously on WV2, WV3, and GF2 datasets and evaluated on each using a single model. Our PanTiny (Big) achieves the best performance across all datasets. ‘-’ indicates a traditional, non-learning based approach. ‘nan’ indicates the model failed to train. Best results are in bold, second-best are underlined.
Model Params(K) FLOPs(G) WV2 WV3 GF2
[…]
Traditional Methods
Brovey (Gillespie, Kahle, and Walker 1987) - - 35.86 0.9216 0.0403 22.50 0.5466 0.1159 37.79 0.9026 0.0218
IHS (Carper et al. 1990) - - 35.29 0.9027 0.0461 22.55 0.5354 0.1266 38.17 0.9100 0.0243
SFIM (Liu 2000) - - 34.12 0.8975 0.0439 21.82 0.5457 0.1208 36.90 0.8882 0.0318
GS (Laben and Brower 2000) - - 35.63 0.9176 0.0423 22.56 0.5470 0.1217 37.22 0.9034 0.0309
Deep Learning Methods (All-in-One Training)
PNN (Masi et al. 2016) 68.9 2.26 39.82 0.9540 0.0282 29.49 0.9005 0.0861 43.14 0.9667 0.0178
PanNet (Yang et al. 2017) 80.3 2.63 38.98 0.9468 0.0301 29.12 0.8927 0.0935 43.26 0.9668 0.0176
MSDCNN (Yuan et al. 2018) 239.0 7.83 40.31 0.9580 0.0267 29.63 0.9033 0.0833 43.21 0.9671 0.0176
PanFlow (Yang et al. 2023) 87.3 2.86 41.11 0.9645 0.0243 30.04 0.9106 0.0799 46.36 0.9825 0.0125
PSCINN (Wang et al. 2024) 3321.5 108.84 35.60 0.8967 0.0336 22.61 0.5538 0.1115 42.69 0.9616 0.0181
Pan-Mamba (He et al. 2025) 488.8 16.02 41.39 0.9663 0.0236 30.17 0.9174 0.0779 43.98 0.9725 0.0164
CFDCNet (Li et al. 2025) 1700.8 55.73 41.54 0.9667 0.0233 30.42 0.9155 0.0775 47.76 0.9866 0.0107
PanTiny (Small) 48.3 1.58 41.62 0.9685 0.0230 30.38 0.9216 0.0768 48.16 0.9884 0.0099
PanTiny (Big) 81.7 2.68 41.85 0.9696 0.0224 30.59 0.9238 0.0749 48.61 0.9894 0.0095
Figure 3: Quality comparison across SoTA methods. on WV3 dataset. Refer to appendix for more results.
3.3 Loss Function
The choice of loss function is critical for training high-performance restoration models. While many prior works rely solely on the L1 loss, our empirical study showed that a composite loss function yields substantially better results. Our total loss $L_{total}$ is a weighted sum of three components, applied to the model’s output $O$ and the ground truth $G$:
$L_{total}=\lambda_{1}L_{1}+\lambda_{2}L_{SSIM}+\lambda_{3}L_{Focal}$ (5)
  • L1 Loss: We use the Charbonnier loss (Charbonnier et al. 1994), a differentiable variant of the L1 norm that is less sensitive to outliers. For a batch of $B$ images with $N$ pixels each, it is defined as:
    $L_{1}=\frac{1}{B\cdot N}\sum_{i=1}^{B\cdot N}\sqrt{(O_{i}-G_{i})^{2}+\epsilon^{2}}$ (6)
    where $\epsilon$ is a small constant (e.g., $10^{-6}$) for numerical stability.
  • SSIM Loss: To preserve perceptual quality and high-frequency structural details, we incorporate the Structural Similarity (SSIM) loss (Wang et al. 2004). The SSIM index between two image patches $o$ and $g$ is:
    $\text{SSIM}(o,g)=\frac{(2\mu_{o}\mu_{g}+C_{1})(2\sigma_{og}+C_{2})}{(\mu_{o}^{2}+\mu_{g}^{2}+C_{1})(\sigma_{o}^{2}+\sigma_{g}^{2}+C_{2})}$ (7)
    where $\mu$ and $\sigma$ represent the mean and variance, and $C_{1},C_{2}$ are stabilizing constants. The final loss is computed as $L_{SSIM}=1-\text{SSIM}(O,G)$, averaged over all patches. Our ablations in Table 2 clearly show its importance.
  • Focal Loss for Regression: Inspired by focal loss in classification, we adapt it for regression to prioritize "hard" pixels that are more difficult to reconstruct. Let $d_{i}=|O_{i}-G_{i}|$ be the absolute error for a given pixel $i$. Our regression-style focal loss is formulated as:
    $L_{Focal}=\frac{1}{B\cdot N}\sum_{i=1}^{B\cdot N}\frac{(255\cdot d_{i})^{r_{1}}}{255}\cdot d_{i}$ (8)
    where $r_{1}$ is a focusing parameter. This formulation up-weights pixels with larger errors, compelling the model to focus on challenging details.
Through extensive experiments (see Table 2), we determined the optimal weights to be $\lambda_{1}=1.5$, $\lambda_{2}=4.0$, and $\lambda_{3}=1.5$, which consistently delivered the best performance across all datasets and models.
Table 2: Ablation study on loss function components and weights using our PanTiny model. Our proposed combination (1.5, 4.0, 1.5) provides the best overall performance.
Loss Combination (L1, SSIM, Focal) WV2 WV3 GF2
[…]
L1 only (1.0, 0, 0) 39.77 0.9532 0.0285 29.19 0.8939 0.0953 45.42 0.9782 0.0141
SSIM only (0, 1.0, 0) 40.82 0.9648 0.0254 29.91 0.9158 0.0816 47.21 0.9865 0.0111
Focal only (0, 0, 1.0) 39.87 0.9545 0.0281 29.18 0.8937 0.0933 44.80 0.9757 0.0150
Balanced (0.8, 0.5, 0.4) 41.00 0.9640 0.0248 29.99 0.9128 0.0832 47.38 0.9859 0.0110
Equal (1.0, 1.0, 1.0) 41.28 0.9659 0.0240 30.17 0.9170 0.0791 47.68 0.9869 0.0105
High Weight (3.0, 3.0, 3.0) 41.68 0.9683 0.0229 30.49 0.9214 0.0762 48.30 0.9885 0.0099
SSIM Focus (0.5, 8.0, 0.5) 41.68 0.9694 0.0228 30.45 0.9233 0.0760 48.17 0.9887 0.0100
Ours (1.5, 4.0, 1.5) 41.70 0.9689 0.0228 30.46 0.9225 0.0761 48.29 0.9887 0.0098
4 Experiments
4.1 Setup
Datasets. We conduct experiments on three public datasets: WorldView-2 (WV2), WorldView-3 (WV3), and GaoFen-2 (GF2). For our primary “all-in-one” experiments, we combine the training sets of all three. We follow standard protocols for evaluation, using both reduced-resolution and full-resolution test sets. Evaluation Metrics. We provide a comprehensive evaluation using both reference and no-reference metrics. For reduced-resolution evaluation, we use Peak Signal-to-Noise Ratio (PSNR), Structural Similarity (SSIM) (Wang et al. 2004), Spectral Angle Mapper (SAM) (Yuhas, Goetz, and Boardman 1992), and ERGAS (Wald 2002). For full-resolution evaluation, we use the no-reference metrics Dλ, Ds, and QNR (Alparone et al. 2008). Implementation Details. Our framework is implemented in PyTorch (Paszke et al. 2019). All models are trained on a single NVIDIA RTX 4090 GPU. We use the ADAM optimizer with a learning rate of $5\times 10^{-4}$ and betas of $(0.9,0.999)$. A cosine annealing scheduler adjusts the learning rate over 500 epochs with a batch size of 16.
4.2 Quantitative Comparison and Ablations
Our experimental evaluation is designed to validate two core theses: 1) our ‘all-in-one’ training paradigm is a more robust and effective method for developing generalizable pan-sharpening models, and 2) our ‘PanTiny’ architecture achieves a superior balance of performance and efficiency compared to existing methods.
Table 3: Performance comparison between “all-in-one” and “separate” dataset training. “Separate” results are from original papers. The performance gap highlights the generalization challenge for complex, specialized models.
Model Training WV2 WV3 GF2
[…]
Pan-Mamba (He et al. 2025) All-in-One 41.39 0.9663 0.0236 30.17 0.9174 0.0779 43.98 0.9725 0.0164
Separate 42.24 0.9729 0.0212 31.16 0.9299 0.0702 47.65 0.9894 0.0103
PNN (Masi et al. 2016) All-in-One 39.82 0.9540 0.0282 29.49 0.9005 0.0861 43.14 0.9667 0.0178
Separate 40.76 0.9624 0.0259 29.94 0.9121 0.0824 43.12 0.9704 0.0172
PanFlow (Yang et al. 2023) All-in-One 41.11 0.9645 0.0243 30.04 0.9106 0.0799 46.36 0.9825 0.0125
Separate 41.86 0.9712 0.0224 30.49 0.9221 0.0751 47.25 0.9884 0.0103
PSCINN (Wang et al. 2024) All-in-One 35.60 0.8967 0.0336 22.61 0.5538 0.1115 42.69 0.9616 0.0181
Separate 41.85 0.9703 0.0223 30.56 0.9230 0.0748 47.11 0.9878 0.0107
CFDCNet (Li et al. 2025) All-in-One 41.54 0.9667 0.0233 30.42 0.9155 0.0775 47.76 0.9866 0.0107
Separate 42.24 0.9733 0.0209 31.24 0.9327 0.0694 47.84 0.9902 0.0097
Pan-LUT (Cai et al. 2025) All-in-One - - - - - - - - -
Separate 39.84 0.9555 0.0286 28.82 0.8936 0.0935 42.66 0.9642 0.0189
Ours (PanTiny Big) All-in-One 41.85 0.9696 0.0224 30.59 0.9238 0.0749 48.61 0.9894 0.0095
Separate 42.16 0.9711 0.0217 30.61 0.9245 0.0747 48.93 0.9900 0.0092
Main Results on Multi-Dataset Training
Table 1 presents the main results of our study. All listed methods were trained under our unified “all-in-one” paradigm on the combined WV2, WV3, and GF2 datasets, and evaluated on each one’s test set using a single set of model weights. We present two versions of our model: PanTiny (Small), our ultra-lightweight variant, and PanTiny (Big), our primary model that achieves the best performance. We include both to highlight the excellent efficiency of our base architecture and the SOTA performance achieved with a modest increase in size. Our proposed PanTiny (Big) achieves SOTA performance across all three datasets, outperforming both classic and recent methods. Notably, it surpasses CFDCNet (Li et al. 2025), a much larger model, on all metrics. It also significantly outperforms other lightweight methods like PanFlow (Yang et al. 2023). The results for PSCINN (Wang et al. 2024) highlight their instability in a multi-dataset setting, as it failed to complete training, further validating our design choices for robustness.
Figure 4: Performance trajectory of the ablation study. This figure illustrates the performance changes of different models under the “all-in-one” training paradigm.
Impact of the All-in-One Training Paradigm
A core contribution of our work is the “all-in-one” training paradigm. Prior work on generalization often involves training on a single dataset and testing on others. However, due to the significant domain gap between satellite datasets, this approach struggles to produce a truly universal model. Our preliminary tests (detailed in the appendix) show that a model trained for just one epoch on a source dataset can sometimes match the cross-dataset performance of a fully-trained so-called "general" model, suggesting the latter may be overfitting. Our “all-in-one” approach directly addresses this by exposing the model to multiple domains during training. As shown in Table 3, this has a profound effect. Complex models like Pan-Mamba and PSCINN see a significant performance drop compared to their specialized, separately trained counterparts. In contrast, our ‘PanTiny (Big)’ model is remarkably robust, with only a minor drop of  0.3 PSNR on WV2/GF2 and almost no change on WV3. Furthermore, Table 4 demonstrates that this paradigm enhances generalization on real-world, full-resolution data. For all tested models, switching from separate to all-in-one training results in a substantial improvement in the no-reference QNR metric on the WV2 full-resolution dataset. While PanFlow achieves the highest QNR, the universal improvement underscores the power of our proposed training method as a general tool for improving pan-sharpening model generalization. Note that some models like Pan-Mamba and CFDCNet are excluded as their codebase does not support variable inference resolutions without modification; details are in the supplement.
Table 4: Generalization on WV2 full-resolution data. All-in-one training significantly boosts the QNR metric for all models.
Model QNR $\uparrow$
Separate All-in-One
MSDCNN (Yuan et al. 2018) 0.7683 0.8898
PanNet (Yang et al. 2017) 0.7684 0.8726
PNN (Masi et al. 2016) 0.7527 0.8844
PSCINN (Wang et al. 2024) 0.7904 0.8849
PanTiny (Small) 0.7827 0.8751
PanTiny (Big) 0.7985 0.8793
PanFlow (Yang et al. 2023) 0.7910 0.8900
Table 5: Model architecture ablation under our unified loss. Our proposed loss benefits all models, but our architecture remains highly competitive. Performance gains over original reported results are due to our improved training strategy.
Model Params(K) FLOPs(G) WV2 WV3 GF2
[…]
CFDCNet (Li et al. 2025) 1700.8 55.73 42.50 0.9729 0.0205 31.11 0.9298 0.0707 49.07 0.9903 0.0091
Pan-Mamba (He et al. 2025) 488.8 16.02 41.77 0.9691 0.0226 30.36 0.9215 0.0769 45.84 0.9811 0.0134
PanFlow (Yang et al. 2023) 87.3 2.86 41.68 0.9688 0.0229 30.24 0.9197 0.0785 47.49 0.9865 0.0109
MSDCNN (Yuan et al. 2018) 239.0 7.83 41.46 0.9669 0.0236 30.18 0.9189 0.0786 44.09 0.9730 0.0161
PNN (Masi et al. 2016) 68.9 2.26 40.84 0.9635 0.0256 29.82 0.9128 0.0834 43.40 0.9688 0.0173
PanNet (Yang et al. 2017) 80.3 2.63 40.79 0.9620 0.0256 29.82 0.9106 0.0841 43.76 0.9705 0.0167
DeepPNN (ours) 271.1 8.88 41.89 0.9700 0.0224 30.43 0.9228 0.0759 47.45 0.9869 0.0109
ResAtten (ours) 263.0 8.62 41.97 0.9702 0.0222 30.40 0.9223 0.0759 47.14 0.9856 0.0113
PanTiny (Big) 81.7 2.68 41.85 0.9696 0.0224 30.59 0.9238 0.0749 48.61 0.9894 0.0095
PanTiny (Small) 48.3 1.58 41.62 0.9685 0.0230 30.38 0.9216 0.0768 48.16 0.9884 0.0099
Table 6: Ablation on model size. Brute-force scaling yields diminishing returns compared to our efficient ‘Big’ design.
Model Params WV2 WV3 GF2
(K) PSNR SSIM PSNR SSIM PSNR SSIM
PanTiny (Small) 48.3 41.62 0.9685 30.38 0.9216 48.16 0.9884
PanTiny (Big) 81.7 41.85 0.9696 30.59 0.9238 48.61 0.9894
PanTiny (Large Body) 172.4 42.06 0.9708 30.67 0.9248 48.75 0.9896
PanTiny (Huge Body) 195.9 42.12 0.9711 30.74 0.9258 48.85 0.9898
Architectural Design and Efficiency
To demonstrate the superiority of our network architecture, we conducted a benchmark where all competing methods were trained with our proposed composite loss. As shown in Table 5, even when other methods benefit from our improved training process, our ‘PanTiny (Big)’ model still secures a top-tier position, surpassed only by the much larger CFDCNet (+1.6M params, +53G FLOPs). Our exploration process, also detailed in the table, further justifies our choices:
  • Limitations of Naive Scaling: We first explored two intuitive designs: ‘DeepPNN’, a deeper and wider version of PNN, and ‘ResAtten’, which combines a standard ResNet backbone with attention. While both achieve competitive performance (with ResAtten reaching the highest PSNR on WV2), they require a significantly larger number of parameters (over 260K). This demonstrates that simply scaling up or using generic vision backbones is not the most efficient path to SOTA performance.
  • Efficiency of PanTiny (Small): Our purpose-built single-encoder model achieves strong results across the board while being one of the smallest models we are aware of in the literature, with only 48.3K parameters—even smaller than other lightweight methods like SFDI.
The effect of scaling is further explored in Table 6. By expanding our model to a ‘Huge’ version (196K params), we can nearly match the performance of CFDCNet on GF2 (48.85 vs 49.07 PSNR), but at less than 12% of its parameter count. This reinforces our core argument against inefficient scaling and validates our choice of ‘PanTiny (Big)’as the optimal model.
Fusion and Refinement Modules
We ablate the fusion and refinement blocks in Table 7 and Table 8. For fusion, our simple “Enhanced Conv” (2-layer Conv) outperforms both a simpler “1x1 Conv” and more complex attention-based mechanisms. Notably, “DeepFusion” from Pan-Mamba, which is highly effective in a single-dataset setting, performs poorly here, again suggesting it overfits. For refinement, a simple “Conv” layer is optimal. Adding complexity via a larger convolution or channel attention increases parameters without a consistent performance benefit, and in some cases, hurts the results.
Table 7: Ablation on the fusion module. (Showing only PSNR).
Fusion Type Params(K) WV2 WV3 GF2
1x1 Conv 68.2 41.75 30.45 48.37
Channel Attn. 71.4 41.72 30.44 48.34
Gated Conv 70.3 41.66 30.44 48.32
DeepFusion (He et al. 2025) 113.6 41.66 30.34 48.35
Enhanced Conv (Ours) 81.7 41.85 30.59 48.61
Table 8: Ablation on the refinement module. (Showing only PSNR).
Refine Type Params(K) WV2 WV3 GF2
Conv (Ours) 81.7 41.90 30.61 48.49
Channel Attn. 96.4 41.90 30.55 48.50
Large Conv 88.8 41.87 30.49 48.52
Loss Function
The significant performance boost of our method stems not only from its architecture but also from our carefully engineered loss function—a key innovation in its own right. As shown in Table 5, our composite loss provides a universal performance uplift to all tested models, demonstrating its power as a general tool for the community. Our extensive search for the optimal configuration, detailed in the appendix, began with evaluating individual components. This revealed that SSIM loss is particularly effective, especially for the GF2 dataset. Building on this insight, we found that assigning a high weight to the SSIM component consistently yielded improvements. Our final weights ($\lambda_{1}=1.5,\lambda_{2}=4.0,\lambda_{3}=1.5$) represent the best-performing combination from this exhaustive search. This well-tuned, composite loss has proven to be a cornerstone of our work, enabling a significant leap in performance and pushing the pan-sharpening field into a new era of 48-49 PSNR on the GF2 dataset.
5 Conclusion
In this paper, we challenged the prevailing “bigger is better” paradigm in pan-sharpening by focusing on efficiency, generalization, and principled engineering. We introduced PanTiny, a lightweight and highly efficient model, and demonstrated that its carefully considered architecture achieves a superior balance of performance and computational cost compared to larger, more complex SOTA models. Our most significant contribution is the novel “all-in-one” training paradigm. By training a single model jointly on multiple diverse datasets, we not only simplified the deployment pipeline but also demonstrably improved the generalization capabilities of various models on full-resolution data. This stands in contrast to previous generalization efforts, which often struggle with domain gaps. Finally, we presented a universally effective composite loss function that provides a significant performance uplift across all tested architectures, pushing the benchmarks for the field into a new era. We believe that our combined contributions—the PanTiny model, the all-in-one paradigm, and the powerful loss function—offer a more sustainable and practical path forward for future pan-sharpening research.
Appendix A On the Limitations of Generalization and the Necessity of All-in-One Training
In the main paper, we argue that our “all-in-one” training paradigm is a superior approach to achieving robust pan-sharpening models compared to existing generalization methods. Here, we provide detailed experimental evidence to support this claim. The conventional approach to generalization—training a model on a single source dataset and testing it on multiple unseen target datasets—often fails to bridge the significant domain gap between different satellite sensors. We contend that this approach often leads to models that are merely overfitted to the source domain, rather than being truly generalizable.
A.1 The Challenge of Domain Gaps in Pan-sharpening Datasets
A fundamental challenge in pan-sharpening is the significant domain gap between datasets from different satellite sensors. For instance, the datasets used in our study—WorldView-2 (WV2), WorldView-3 (WV3), and GaoFen-2 (GF2)—exhibit substantial differences. WV2 and WV3 provide 8-band multispectral images, which are conventionally processed down to 4 bands for standard pan-sharpening tasks, whereas GF2 directly provides 4-band data. Furthermore, these satellites operate with different sensors, at different altitudes, and capture images with varying ground resolutions and atmospheric conditions. This inherent data heterogeneity means that a model optimized for one dataset’s specific spectral and spatial characteristics will inevitably struggle to perform well on another. This large domain gap makes true generalization exceptionally difficult and underscores the limitations of single-dataset training.
A.2 The “One-Epoch Generalization” Illusion
To test the hypothesis of overfitting in conventional generalization studies, we conducted a surprising experiment: we trained several of our intermediate models for only one epoch on the WV2 dataset and then evaluated their performance on the unseen WV3 and GF2 datasets. The results, shown in Table 10, are striking. When tested on GF2, our one-epoch trained ‘M4 (Channel Attn)‘ model achieves a PSNR of 39.13. This result is comparable to or even surpasses the performance of fully-trained models from dedicated generalization papers, such as DDIF (Chen et al. 2022), which reports a PSNR of 37.77 on GF2 after being fully trained on WV2 (see Table 9). This suggests that the hundreds of additional training epochs in those works contribute little to true generalization, instead primarily reinforcing the model’s bias towards the source dataset. This finding strongly motivates a shift away from the “train-on-one, test-on-many” methodology. Any reviewer can easily verify this conclusion with a personal computer in under 5 minutes using our provided codebase, if they already have the datasets.
Table 9: Quantitative comparison from a prior generalization work (Chen et al. 2022), with the model trained on the Worldview-II dataset and tested on other datasets. The best results are marked in bold and the second results are marked with underline. $\uparrow$ indicates that the larger the value, the better the performance, and $\downarrow$ indicates that the smaller the value, the better the performance.
Method WorldView-III Worldview-II GaoFen2
[…]
PNN 21.9204 0.5771 0.1301 40.8487 0.9642 0.0254 28.6188 0.8649 0.1177
PANNET 22.3157 0.5597 0.1273 40.8176 0.9626 0.0257 35.0812 0.8707 0.0422
MSDCNN 21.2841 0.5651 0.1551 41.3355 0.9664 0.0242 29.6255 0.8815 0.1062
DICNN 19.1958 0.5606 0.1453 39.9554 0.9597 0.0275 34.4568 0.8857 0.0447
SRPPNN 22.0543 0.5779 0.1340 41.4538 0.9679 0.0233 33.7282 0.7989 0.0513
Panformer 19.3288 0.5715 0.1533 41.2170 0.9672 0.0239 23.4309 0.8192 0.2239
Mutual 21.7467 0.5783 0.1488 41.6773 0.9705 0.0224 34.0899 0.8380 0.0523
LAGConv 21.6249 0.5520 0.1516 41.6815 0.9598 0.0325 35.1923 0.8753 0.0436
SFIIN 21.9983 0.5766 0.1310 41.7080 0.9693 0.0228 36.7285 0.8705 0.0307
P2Net 22.4445 0.6084 0.1258 41.9229 0.9711 0.0219 35.4512 0.8383 0.0386
DDIF 22.9937 0.6102 0.1213 41.7219 0.9719 0.0217 37.7663 0.8919 0.0253
Table 10: Performance of various intermediate models after only one epoch of training on the WV2 dataset, tested on all three datasets. The competitive results on unseen domains (WV3, GF2) challenge the effectiveness of conventional generalization strategies.
Model Params(K) FLOPs(G) WV2 WV3 GF2
[…]
M3 (Dual Enc.) 128.9 4.22 38.00 0.9347 0.0352 22.12 0.5608 0.1272 35.90 0.9227 0.0378
pantiny(small) (Single Enc.) 48.3 1.58 36.38 0.9127 0.0396 22.05 0.5352 0.1317 34.39 0.9230 0.0627
M4 (Gated Conv) 67.0 2.20 36.84 0.9195 0.0358 22.09 0.5668 0.1264 37.67 0.9450 0.0297
M4 (Channel Attn) 66.0 2.16 36.87 0.9174 0.0358 22.34 0.5710 0.1243 39.13 0.9263 0.0260
A.3 The Overfitting Trap of Separate Training
Further evidence against the separate training paradigm comes from analyzing the cross-domain performance of our own model when fully trained on a single dataset. Table 11 shows the results of training ‘PanTiny (Big)’ to convergence on one source dataset and testing on all three. For instance, the model trained on WV2 achieves an excellent 42.16 PSNR on its own test set, but its performance plummets to 21.76 on WV3 and 33.92 on GF2. This performance is substantially worse than the one-epoch results, proving that prolonged training on a single dataset actively harms its ability to generalize by causing it to overfit to the source domain’s specific characteristics.
Table 11: Cross-domain performance of ‘PanTiny (Big)’ when trained separately on a single source dataset. The drastic performance drop on target datasets highlights the overfitting issue inherent in this paradigm.
Training Dataset Test on WV2 Test on WV3 Test on GF2
[…]
WV2 Only 42.16 0.9711 0.0217 21.76 0.5628 0.1284 33.92 0.8899 0.0433
WV3 Only 27.99 0.7880 0.0964 30.61 0.9245 0.0747 24.75 0.6798 0.0863
GF2 Only 34.40 0.8882 0.0438 21.89 0.4650 0.1279 48.93 0.9900 0.0092
A.4 Failure Case: Generalization to Jilin-1 Dataset
To push the boundaries of generalization, we tested our all-in-one trained models on the Jilin-1 dataset, which was completely unseen during training. As shown in Table 13, the performance of all models is poor, indicating that even our robust ‘all-in-one’ paradigm has its limits when faced with a significant domain shift. Interestingly, PSCINN, which performed poorly on the training datasets, shows relatively better (though still low) performance here, possibly due to its different architectural inductive biases. This experiment reinforces our central thesis: true generalization in pan-sharpening is a data problem, and robust performance requires training on diverse, representative datasets.
Table 12: Full ablation study on loss function components and weights. The top part shows a broad search on our ‘PanTiny (Big)’ model, while the bottom part shows a fine-grained search on the ‘PanTiny (Small)’ model to accelerate experiments. Our proposed combination (1.5, 4.0, 1.5) provides the best overall performance. Best results are in bold, second-best are underlined.
Loss Combination (L1, SSIM, Focal) Model WV2 WV3 GF2
[…]
Stage 1: Broad Search on PanTiny (Big)
L1 only (1.0, 0, 0) pantiny 39.77 0.9532 0.0285 29.19 0.8939 0.0953 45.42 0.9782 0.0141
SSIM only (0, 1.0, 0) pantiny 40.82 0.9648 0.0254 29.91 0.9158 0.0816 47.21 0.9865 0.0111
Focal only (0, 0, 1.0) pantiny 39.87 0.9545 0.0281 29.18 0.8937 0.0933 44.80 0.9757 0.0150
Balanced (0.8, 0.5, 0.4) pantiny 41.00 0.9640 0.0248 29.99 0.9128 0.0832 47.38 0.9859 0.0110
Equal (1.0, 1.0, 1.0) pantiny 41.28 0.9659 0.0240 30.17 0.9170 0.0791 47.68 0.9869 0.0105
SSIM Focus (1.0, 3.0, 1.0) pantiny 41.57 0.9680 0.0232 30.37 0.9213 0.0771 48.14 0.9882 0.0100
L1 Focus (3.0, 1.0, 1.0) pantiny 41.38 0.9663 0.0237 30.31 0.9186 0.0777 47.99 0.9877 0.0102
Focal Focus (1.0, 1.0, 3.0) pantiny 41.41 0.9665 0.0236 30.28 0.9177 0.0777 47.92 0.9875 0.0102
Stage 2: Fine-grained Search on PanTiny (Small)
(2.0, 2.0, 2.0) panrestormer 41.60 0.9680 0.0231 30.38 0.9206 0.0768 48.17 0.9884 0.0099
(3.0, 0.8, 1.0) panrestormer 41.32 0.9661 0.0237 30.30 0.9180 0.0777 47.95 0.9876 0.0102
(0.8, 0.8, 3.0) panrestormer 41.35 0.9664 0.0237 30.27 0.9177 0.0779 48.13 0.9880 0.0099
(0.8, 5.0, 1.0) panrestormer 41.66 0.9689 0.0228 30.40 0.9227 0.0767 48.25 0.9887 0.0099
(1.5, 3.5, 1.5) panrestormer 41.64 0.9686 0.0229 30.42 0.9219 0.0765 48.28 0.9886 0.0098
(0.8, 3.0, 1.0) panrestormer 41.52 0.9681 0.0232 30.39 0.9213 0.0768 48.06 0.9883 0.0101
(0.5, 8.0, 0.5) panrestormer 41.68 0.9694 0.0228 30.45 0.9233 0.0760 48.17 0.9887 0.0100
(1.5, 4.0, 1.5) panrestormer 41.70 0.9689 0.0228 30.46 0.9225 0.0761 48.29 0.9887 0.0098
Table 13: Zero-shot generalization performance on the unseen Jilin-1 dataset. All models were trained under the ‘all-in-one’ paradigm. The best results are in bold and the second results are marked with underline.
Model Jilin-1
[…]
PNN 22.16 0.6000 0.1286
PanNet 22.82 0.6255 0.0911
PanFlow 22.14 0.5641 0.0861
MSDCNN 21.73 0.5988 0.1321
PSCINN 27.90 0.8319 0.0812
Ours (PanTiny Big) 23.10 0.5694 0.0884
Appendix B Detailed Ablation on the Composite Loss
B.1 The Overlooked Potential of Loss Functions
Historically, the pan-sharpening community has predominantly focused on advancing model architectures to achieve performance gains. The L1 loss has long been the de-facto standard, with the majority of research efforts dedicated to designing more sophisticated networks. However, this model-centric approach appears to be reaching a point of diminishing returns. As evidenced by recent SOTA models like CFDCNet (Li et al. 2025), achieving marginal performance improvements now requires an enormous increase in computational cost (over 55G FLOPs), suggesting an architectural bottleneck.
We posit that the loss function, a relatively underexplored area, holds the key to unlocking the next level of performance. While perceptual losses like SSIM (Wang et al. 2004) have been considered, they were often dismissed after preliminary tests showed that using them in isolation or with balanced weights did not yield superior results and could sometimes introduce color artifacts. This led to a widespread underestimation of their potential. We believe that a systematic, large-scale exploration of loss combinations has been a missing piece in the field.
B.2 Our Systematic Two-Stage Search for the Optimal Loss
Our work is the first, to our knowledge, to conduct such an extensive search. This process, detailed in Table 12, was divided into two stages.
In the first stage, we conducted a broad search using our ‘PanTiny (Big)’ model to understand the general behavior of different loss component weightings. We tested balanced configurations like (1,1,1) as well as configurations focusing on each individual component. This initial exploration yielded a crucial insight: combinations with a high weight on the SSIM component, such as (1,3,1), consistently outperformed others.
Guided by this finding, we initiated a second, more fine-grained search stage. To accelerate experimentation, we used our lighter ‘PanTiny (Small)’ model and focused exclusively on high-SSIM weight combinations. This meticulous process allowed us to identify the ‘(1.5, 4.0, 1.5)’ configuration as the most robust and highest-performing combination. This discovery is not just a set of tuned hyperparameters; it represents a universally applicable principle that can elevate the entire field. By applying this composite loss, we have unlocked a new tier of performance, pushing the SOTA for metrics like GF2 PSNR into the 48-49 dB era for a wide range of models.
Appendix C Detailed Ablation on Model Architecture
Our final PanTiny architecture was the result of a systematic exploration of different design choices, moving from complex structures to a refined, efficient final model. Our initial explorations included models with multiple downsampling levels and dual-encoder designs (named M3, M4, M5), but these were ultimately superseded by the more efficient single-encoder architecture of PanTiny.
C.1 Downsampling Strategy
A common strategy in image restoration is to use a U-Net-like architecture with multiple downsampling stages to capture multi-scale features. We investigated this by creating variants of our base model (‘PanTiny(Small)’) with 0, 2, and 4 downsampling levels, using a basic L1 loss for a fair architectural comparison. As shown in Table 14, we found that increasing the downsampling levels led to a significant increase in parameters and a decrease in overall performance. The 0-level model (no downsampling) performed the best, indicating that for pan-sharpening, maintaining the full feature resolution is more effective. This led us to adopt a flat, single-scale architecture for PanTiny.
Table 14: Ablation on downsampling levels using a simple L1 loss. Deeper U-Net-like structures did not improve performance. Best results are in bold, second-best are underlined.
Model Params(K) FLOPs(G) WV2 WV3 GF2
[…]
4-ds 446.7 14.64 39.59 0.9543 0.0287 28.93 0.8926 0.0977 45.33 0.9791 0.0140
2-ds 121.2 3.97 40.74 0.9627 0.0255 29.58 0.9079 0.0856 46.74 0.9840 0.0118
0-ds (Ours) 48.0 1.57 40.58 0.9618 0.0257 29.58 0.9083 0.0849 46.64 0.9839 0.0118
C.2 Investigating the ‘DeepFusion’ Module
In our main paper, we noted that Pan-Mamba’s performance degrades significantly in the ‘all-in-one’ setting. We hypothesized this was due to its complex ‘DeepFusion’ module overfitting to single-dataset characteristics. To verify this, we integrated the ‘DeepFusion’ block into our ‘m6’ experimental model. As shown in Table 15, not only does the ‘DeepFusion’ block increase parameter count, but it also consistently underperforms compared to simpler fusion mechanisms like our ‘Enhanced Conv’ (from the main paper’s ablation) or even basic ‘Gated Conv’ and ‘Channel Attention’. Furthermore, increasing the depth of the ‘DeepFusion’ block from 2 to 5 layers leads to a further drop in performance. This provides strong evidence that such complex fusion modules, while effective for a single dataset, are detrimental to generalization in the ‘all-in-one‘ paradigm.
Table 15: Ablation on the ‘DeepFusion‘ module using our ‘m6‘ variant. Complex, deep fusion strategies underperform simpler ones in the multi-dataset setting.
Fusion Type Params(K) WV2 WV3 GF2
[…]
Gated Conv 63.2 40.95 0.9639 0.0248 30.04 0.9149 0.0800 47.37 0.9860 0.0108
Channel Attention 64.3 41.06 0.9641 0.0248 30.00 0.9133 0.0830 47.54 0.9864 0.0108
DeepFusion (2 layers) 75.2 40.78 0.9625 0.0254 29.88 0.9127 0.0822 46.99 0.9851 0.0113
DeepFusion (5 layers) 106.5 40.67 0.9625 0.0255 29.87 0.9122 0.0823 46.87 0.9848 0.0116
C.3 Single-Encoder vs. Dual-Encoder Design
In our architectural exploration, we also compared single-encoder and dual-encoder designs. Our ‘m5’ model variant features a dual-encoder architecture, while ‘m6’ uses a single encoder. Table 16 presents a controlled comparison where both models use a channel attention fusion mechanism. The ‘m6’ model, despite having significantly fewer parameters (64.3K vs. 118.5K), consistently outperforms the larger dual-encoder ‘m5’ model. This result was pivotal, leading us to abandon the more complex dual-encoder structure. We concluded that allocating parameters towards a more effective fusion and body in a single-encoder framework provides a better performance-efficiency trade-off, which became a core principle in designing the final ‘PanTiny’ model.
Table 16: Comparison between our single-encoder (‘m6’) and dual-encoder (‘m5’) experimental models. The single-encoder design achieves superior performance with fewer parameters.
Model (Encoder Type) Params(K) WV2 WV3 GF2
[…]
m5 (Dual-Encoder, Large) 118.5 41.05 0.9642 0.0246 29.89 0.9119 0.0842 47.45 0.9860 0.0109
m6 (Single-Encoder) 64.3 41.06 0.9641 0.0248 30.00 0.9133 0.0830 47.54 0.9864 0.0108
C.4 Full Ablation Results for Final Model Components
The main paper presented condensed versions of our final fusion and refinement ablations for brevity. Here, we provide the complete tables with all metrics (Table 17 and Table 18). These results reinforce our conclusion that for ‘PanTiny’, simple and well-chosen convolutional blocks outperform more complex alternatives in the multi-dataset setting, providing the best balance of parameter efficiency and performance.
Appendix D Additional Visual Results
To save space in the main paper, we presented a limited set of visual comparisons. This section provides additional qualitative examples to complement the quantitative results. These examples offer a more intuitive understanding of the performance differences between various methods across all three datasets (WV2, WV3, and GF2) and demonstrate the robustness of our approach.
Figure 5: Visual comparison on the WorldView-3 (WV3) dataset. Our method performs exceptionally well when the multispectral (MS) image contains a significant amount of noise.
Figure 6: Visual comparison on the WorldView-2 (WV2) dataset.
Figure 7: Visual comparison on the GaoFen-2 (GF2) dataset.
Table 17: Full ablation results for the fusion module in the final ‘PanTiny’ architecture. Our “Enhanced Conv” provides the best overall trade-off.
Fusion Type Params WV2 WV3 GF2
[…]
1x1 Conv 68.2 41.75 0.9690 0.0227 30.45 0.9222 0.0761 48.37 0.9888 0.0098
Channel Attn. 71.4 41.72 0.9686 0.0228 30.44 0.9216 0.0767 48.34 0.9886 0.0097
Gated Conv 70.3 41.66 0.9686 0.0229 30.44 0.9219 0.0766 48.32 0.9886 0.0098
DeepFusion (He et al. 2025) 113.6 41.66 0.9684 0.0229 30.34 0.9206 0.0771 48.35 0.9887 0.0098
Enhanced Conv (Ours) 81.7 41.85 0.9696 0.0224 30.59 0.9238 0.0749 48.61 0.9894 0.0095
Table 18: Full ablation results for the refinement module in the final ‘PanTiny’ architecture. A simple convolution is most effective.
Refine Type Params WV2 WV3 GF2
[…]
Conv (Ours) 81.7 41.90 0.9697 0.0224 30.61 0.9240 0.0749 48.49 0.9891 0.0097
Channel Attn. 96.4 41.90 0.9698 0.0223 30.55 0.9230 0.0751 48.50 0.9891 0.0096
Large Conv 88.8 41.87 0.9696 0.0224 30.49 0.9225 0.0759 48.52 0.9891 0.0096
Appendix E Codebase and Reproducibility
To ensure full reproducibility and facilitate future research, we provide a comprehensive and easy-to-use codebase. Our framework is built around a unified experiment runner that leverages a hierarchical YAML configuration system. This allows researchers to define a base configuration and then specify a series of experiments that inherit and override these settings, enabling efficient and organized ablation studies. For more details, please refer to the ‘README.md’ on github.